# ChatDBG: An AI-Powered Debugging Assistant

### v2

@@ Abstract.
Debugging is a critical but challenging task for programmers. This paper proposes ChatDBG, an AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state. It then reports its findings and yields back control to the programmer. Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded roughly 50,000 times.
@@
$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingconference: Under Submission; 2025; Someplace, Somewhere
@@ 1. Introduction
Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task. Programmers must still reason about program behavior to ascertain what went wrong. They must formulate and test hypotheses about program execution, they must read and understand code they may have not written, and they must pore over potentially voluminous information. Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
This paper introduces ChatDBG, the first AI-powered debugger assistant. ChatDBG integrates into and significantly extends the functionality of standard debuggers. ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 (OpenAI, 2023b), enable a debugger to leverage insights and intuition from the vast real-world knowledge embedded in LLMs.
A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’. The question can be as simple as ‘why?’ if a program has crashed or failed an assertion. To answer such queries, ChatDBG orchestrates a conversation with an LLM. A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent while answering the programmer’s queries. Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages. Support for the LLM’s agentic control of the debugger is a key technical contribution.
@@
  • It introduces ChatDBG, the first AI-powered debugger assistant.
@@
  • It introduces the “take the wheel” approach to integrating agentic reasoning in large language models with developer tools.
@@
ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers. ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a natural, conversational, and interactive way.
@@
Figure 1. An example program containing several bugs. It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
@@
Figure 2. A debugging session with ChatDBG. For brevity, we elide repetitive or unimportant parts. ChatDBG catches the assertion failure on line 15 and prompts the user to enter commands (bold and highlighted in gray). Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
@@ 2. Overview
Figure 3. A debugging session demonstrating how ChatDBG incorporates real-world knowledge. After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value. ChatDBG identifies high variance resulting from the small number of trials as the root cause.
@@ 3. Related Work
Table 1. Debugger features and their dates of introduction. Most key features have been around for decades. By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state. (An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
@@
Fault localization seeks to identify the likely location of a defect’s root cause. Several prior studies have investigated the use of machine learning and LLMs for fault localization. Some of the studied techniques apply machine learning to source code features, coverage data, or other static code features to predict faulty lines of code, but they do not utilize dynamic state and run-time information. DeepFL (Li et al., 2019), Grace (Lou et al., 2021) and DeepRL4FL (Li et al., 2021) are examples of such systems. Similarly, LLMAO (Yang et al., 2024) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code. AutoFL (Kang et al., 2024) also utilizes an LLM and enables it to statically retrieve source code and coverage information about the program via function calls. However, the system requires a failing test case as input and does not employ run-time state information.
ChatDBG improves upon these systems by providing an LLM with access to run-time program state and the ability to take control of the underlying debugger. Both of these features enhance the LLM’s ability to provide more accurate and informative feedback to the user. We also note that other fault localization techniques can be used in tandem with ChatDBG to improve results, as suggested by Section 5.1’s utilization of backwards slicing to identify code relevant to a bug in Python notebooks and Section 5.2’s utilization of AddressSanitizer (Serebryany et al., 2012) to provide a better starting point for diagnosing and fixing memory errors in native code.
@@ 4.2. Debugging a Target Program
Figure 4. ChatDBG architecture and command processing algorithm.
@@ 4.3. ChatDBG Architecture Overview
Figure 5. The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
@@ 4.4. Initial Prompts and Enriched Stack Traces
In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error. We illustrate the components of the prompt in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2 as a running example.
@@
In Python, ChatDBG leverages Pdb’s internal data structures to build enriched stack traces. When converting values to suitable string representations, ChatDBG must balance utility with the size of the string produced. For objects, ChatDBG calls the object’s __repr__ method if an appropriate (non-default) version exists. Otherwise, it iterates over the object’s fields and recursively converts their values to strings. Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number. The rest of the elements are just abbreviated with an ellipsis (…). This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses. This strategy balances the need to provide the LLM with sufficient information to diagnose the error with the need to avoid overwhelming it with too much information. In cases with the elided details are important, the LLM can request them via the take the wheel mechanism.
@@
Error. A description of the error causing execution to stop is also extracted from the underlying debugger. When the error is due to an assertion failure, ChatDBG instructs the LLM to assume that the assertion is valid as written so that it will look beyond the assertion for the real problem.
@@ 4.5. Taking the Wheel
ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models (OpenAI, 2023a). This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response. The LLM calls these functions by sending special messages to the client as part of its response stream. The client receives those messages, computes the requested results, and sends them back to the LLM. The initial prompt describes how to use the available functions.
@@
Supporting agentic reasoning over run-time program state via function calls is a key technical innovation of ChatDBG. Without this capability, there would be no effective way to provide the LLM with a detailed view of relevant program state. A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (DBLP:conf/nips/LewisPPPKGKLYR020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval. However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
@@
Table 2. ChatDBG command extensions. ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code. ChatDBG provides access to them via the LLM’s function call API.
@@ 4.6. Navigating the Code
While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions. Trying to include all possibly-relevant source code in the initial prompt would be infeasible and would prevent ChatDBG from scaling to larger codebases. Instead, ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code. These commands are available to the LLM via function calls and are listed in Table 2.
@@ 4.8. Security and Risks
Table 3. Python programs exhibiting a variety of common errors. Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code. Semantic errors reflect failed tests expressed as assertions. Crashes reflect unexpected termination due to any other type of error.
@@ 5. Evaluation
We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs. The former have well-defined expected behavior that enable us to thoroughly and systematically assess ChatDBG. The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes. Our evaluation addresses the following research questions: RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python? RQ2: Which components of ChatDBG contribute to its effectiveness? RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
@@ 5.1. Python
We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table 3. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
Unlike many existing bug benchmarks for Python, these programs are unpublished and thus not in the language model’s training data. In addition, the programs have clear correctness criteria that lead to objective effectiveness metrics in our experiments. The bugs are also representative of mistakes often made in languages like Python. They range from scoping issues, algorithmic errors, and misuse of library functions to subtle misunderstanding of domain knowledge. They include both semantic errors leading to failed tests and crashing errors that terminate execution abruptly. Further, they reflect two important, widely-used modalities for Python programming: non-interactive scripts and interactive computational notebooks. ChatDBG supports debugging in both settings.
Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions. We ran each program ten times under the five configurations in Table 4: Default Stack includes standard stack traces, as generated by ipdb (ipdb, 2007), with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel. Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands. These three configurations all use why? as the initial user text. +Targeted Question asks a question specific to the failure. For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they are intended to represent. For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
@@
Table 4. Configurations used in the Python experiments.
Figure 6. Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
@@
Figure 6 shows the success rate under each configuration. With all features enabled, ChatDBG was successful at identifying and fixing bugs in well over half of the trials. Any time or energy expended by the user manually debugging those cases would be all but eliminated by using ChatDBG.
@@
Figure 7 presents the success rates for each program under each configuration. The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particularly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure. However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
The LLM most heavily used the info, slice (for notebooks), and p (print) debugging commands. That is not surprising, given that they often provide the most direct insight into the execution state and code. The slice command was critically important for notebooks. Without it, success rates rarely improved when the LLM took the wheel.
Figure 7. Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
@@
Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
@@
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people!) to find it. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow ups.

### v3

# ChatDBG: Augmenting Debugging with Large Language Models

@@ Abstract.
Debugging is a critical but challenging task for programmers. This paper proposes ChatDBG, an AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state. It then reports its findings and yields back control to the programmer. By leveraging the real-world knowledge embedded in LLMs, ChatDBG can diagnose issues identifiable only through the use of domain-specific reasoning. Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded more than 65,000 times.
@@
$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009a; Eisenstadt, 1993; Layman et al., 2013). Programmers must still reason about program behavior to ascertain what went wrong. They must formulate and test hypotheses about program execution, they must read and understand code they may have not written, and they must pore over potentially voluminous information. Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
This paper introduces the ChatDBG AI-powered debugger assistant. ChatDBG integrates into and significantly extends the functionality of standard debuggers. ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 (OpenAI, 2023b), enable a debugger to leverage insights and intuition from the vast real-world knowledge embedded in LLMs. This knowledge enables ChatDBG to fix classes of issues that depend on logical thinking and domain-specific reasoning beyond the ability to write and debug programs. For example, Figure 2 illustrates the use of ChatDBG to debug a program by leveraging a knowledge of statistics that cannot be gleaned from the program itself.
A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’. The question can be as simple as ‘why?’ if a program has crashed or failed an assertion. To answer such queries, ChatDBG orchestrates a conversation with an LLM. A key feature of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent (Wooldridge and Jennings, 1995; Gur et al., 2024) while answering the programmer’s queries. Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
@@
  • It introduces ChatDBG, an AI-powered debugger assistant that enables large language models to “take the wheel” and control the debugger via agentic reasoning.
@@
Our evaluation shows that ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers. ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a conversational and interactive way.
@@
Figure 1. An example program containing several bugs (§2). It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
@@
Figure 2. A debugging session with ChatDBG2). For brevity, we elide repetitive or unimportant parts. ChatDBG catches the assertion failure on line 15 and prompts the user to enter commands (bold and highlighted in gray). Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
@@ 2. Overview
Figure 3. A debugging session demonstrating how ChatDBG incorporates real-world knowledge (§2). After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value. ChatDBG identifies high variance resulting from the small number of trials as the root cause.
@@ 3. Related Work
Table 1. Debugger features and their dates of introduction (§3). Most key features have been around for decades. By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state. (An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
@@
Fault localization seeks to identify the likely location of a defect’s root cause. Several prior studies have investigated the use of machine learning and LLMs for fault localization. Some of the studied techniques apply machine learning to source code features, coverage data, or other static code features to predict faulty lines of code, but they do not utilize dynamic state and run-time information. DeepFL (Li et al., 2019), Grace (Lou et al., 2021) and DeepRL4FL (Li et al., 2021) are examples of such systems. Similarly, LLMAO (Yang et al., 2024b) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code. AutoFL (Kang et al., 2024) also utilizes an LLM and enables it to statically retrieve source code and coverage information about the program via function calls. However, the system requires a failing test case as input and does not employ run-time state information.
ChatDBG improves upon these systems by providing an LLM with access to run-time program state and the ability to take control of the underlying debugger. Both features enhance the LLM’s ability to provide more accurate and informative feedback to the user. We also note that other fault localization techniques can be used in tandem with ChatDBG to improve results, as suggested by Section 5.1’s utilization of backwards slicing to identify code relevant to a bug in Python notebooks and Section 5.2’s utilization of AddressSanitizer (Serebryany et al., 2012) to provide a better starting point for diagnosing and fixing memory errors in native code.
@@
3.1. Concurrent work
Several approaches developed concurrently with ChatDBG have also integrated LLMs into automatic program repair or fault localization techniques to enhance the debugging process. Robin (Bajpai et al., 2024) is a chat-based debugging assistant designed to help users diagnose errors more quickly. Both Robin and ChatDBG provide a limited program context to the LLM at the beginning of a conversation. However, Robin has no direct access to any additional context about the program and execution state; the user must manually retrieve and provide these items. Robin’s functionality is therefore roughly equivalent to the Enriched Stack configuration of ChatDBG, detailed in Section 5.1. As Figure 6 shows, ChatDBG achieves a nearly two-fold increase improvement in diagnosing errors versus the Enriched Stack configuration. ChatDBG’s effectiveness generally increases further with targeted questions and follow-up discussions with the user.
AutoCodeRover (Zhang et al., 2024) and SWE-agent (Yang et al., 2024a) are complementary approaches that focus on fault localization and automatic repair, relying exclusively on issue descriptions and source code. ChatDBG additionally leverages run-time information to identify root causes and propose fixes. Section 5 demonstrates the strength of this approach over relying solely on static information. Both AutoCodeRover and SWE-agent perform an evaluation using SWE-bench (Jimenez et al., 2024), which was created to evaluate the efficacy of such static tools; unfortunately, this benchmark suite is not applicable to ChatDBG due to its extensive usage of run-time information.
RepairAgent (Bouzenia et al., 2024) and AutoSD (Kang et al., 2025) are tools that employ LLMs in specific workflows that mimic standard debugging strategies in an attempt to repair pre-identified bugs. While successful in some settings, both tools rely on the user providing a failing test case and the precise location of the bug. By contrast, ChatDBG does not require this information. ChatDBG also enables a more flexible workflow that permits collaboration with the developer in addition to seamless integration into the standard debugging process.
@@ 4.2. Debugging a Target Program
Figure 4. ChatDBG architecture and command processing algorithm (§4.3).
@@ 4.3. ChatDBG Architecture Overview
Figure 5. The initial prompt for the debugging session in Figure 24.4). For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
@@ 4.4. Initial Prompts and Enriched Stack Traces
In addition to including the user’s text, the initial prompt conveys instructions to LLM and the context surrounding the error. We illustrate the components of the prompt in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2 as a running example.
@@
In Python, ChatDBG leverages Pdb’s internal data structures to build enriched stack traces. When converting values to suitable string representations, ChatDBG must balance utility with the size of the string produced. For objects, ChatDBG calls the object’s __repr__ method if an appropriate (non-default) version exists. Otherwise, it iterates over the object’s fields and recursively converts their values to strings. Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number. The rest of the elements are abbreviated with an ellipsis (…). This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses. This strategy balances the need to provide the LLM with sufficient information to diagnose the error with the need to avoid overwhelming it with too much information. In cases with the elided details are important, the LLM can request them via the take the wheel mechanism.
@@
Error. A description of the error causing execution to stop is extracted from the underlying debugger. When the error is due to an assertion failure, ChatDBG instructs the LLM to assume that the assertion is valid as written so that it will look beyond the assertion for the real problem.
@@ 4.5. Taking the Wheel
ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models (OpenAI, 2023a). This agentic approach (Wooldridge and Jennings, 1995; Gur et al., 2024) lets clients register callback functions with the LLM for obtaining additional information while constructing a response. The LLM calls these functions by sending special messages to the client as part of its response stream. The client receives those messages, computes the requested results, and sends them back to the LLM. The initial prompt describes how to use the available functions.
@@
Supporting agentic reasoning over run-time program state via function calls is a key technical innovation of ChatDBG. Without this capability, there would be no effective way to provide the LLM with a detailed view of relevant program state. A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (Lewis et al., 2020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval. However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
@@
Table 2. ChatDBG command extensions (§4.6). ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code. ChatDBG provides access to them via the LLM’s function call API.
@@ 4.6. Navigating the Code
While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions. Trying to include all possibly-relevant source code in the initial prompt would be infeasible and would prevent ChatDBG from scaling to larger codebases. Instead, ChatDBG extends the underlying debuggers with several new commands designed to help the LLM navigate through and understand the target’s code. These commands are available to the LLM via function calls and listed in Table 2.
@@ 4.8. Security and Risks
Table 3. Python programs exhibiting a variety of common errors (§5.1). Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code. Semantic errors reflect failed tests expressed as assertions. Crashes reflect unexpected termination due to any other type of error.
@@ 5. Evaluation
We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs. The former have well-defined expected behavior that enables us to thoroughly and systematically assess ChatDBG. The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes. Our evaluation addresses the following research questions: RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python? RQ2: Which components of ChatDBG contribute to its effectiveness? RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
@@ 5.1. Python
We applied ChatDBG to all bugs in a collection of student labs from two introductory computer science courses; see Table 3. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
Unlike many existing bug benchmarks for Python, these programs are unpublished and thus not in the language model’s training data. In addition, the programs have clear correctness criteria that lead to objective effectiveness metrics in our experiments. The bugs are representative of common mistakes because they were introduced by real humans, rather than synthetically generated. They range from scoping issues, algorithmic errors, and misuse of library functions to subtle misunderstanding of domain knowledge. They include both semantic errors leading to failed tests and crashing errors that terminate execution abruptly. Further, they reflect two important, widely-used modalities for Python programming: non-interactive scripts and interactive computational notebooks. ChatDBG supports debugging in both settings.
Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions. We focus our evaluation on Python and perform an ablation study by progressively enabling ChatDBG features. We ran each program ten times under the five configurations in Table 4: Default Stack includes standard stack traces, as generated by ipdb (ipdb, 2007), with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel. Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands. These three configurations all use why? as the initial user text. +Targeted Question asks a question specific to the failure. For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they intend to represent. For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
@@
Table 4. Configurations used in the Python experiments (§5.1).
Figure 6. Overall ChatDBG success rate for each configuration (§5.1). ChatDBG innovations and user-provided context gradually increase effectiveness.
@@
Figure 6 shows the success rate under each configuration. The simplest configuration, Default Stack, provides functionality roughly equivalent to the user copying and pasting the program stack trace and basic error information into an LLM chat window and requesting a fix. We use this configuration as a baseline for evaluating the impact of ChatDBG’s more advanced configurations. With all features enabled, ChatDBG was successful at identifying and fixing bugs in well over half of the trials. Any time or energy expended by the user manually debugging those cases would be all but eliminated by using ChatDBG.
@@
Figure 7 presents the success rates for each program under each configuration. The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particularly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure. However, enriched stacks coupled with letting the LLM take the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
Using the +Take the Wheel feature, the LLM issues from 0 to 12 debugging commands per run, most commonly calling the info, slice (for notebooks), and p (print) debugging commands. While all of these commands provide useful information about execution state and code, the slice command was critically important for notebooks. Without it, success rates rarely improved when the LLM took the wheel.
Figure 7. Success rate for ChatDBG for each program and configuration (§5.1). Vertical lines show the mean.
@@
Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each (§5.2).
@@
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. We observe that in some cases, enabling additional features in ChatDBG decreases its success rate. We attribute this result to the fact that longer and more complex prompts can occasionally degrade the effectiveness of LLMs (Levy et al., 2024). In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009b, p. 243)) to find it. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow-ups.